Can someone explain how to do this problem?

2. Let f(t, y) — х +у, 0<x< 1, 0 <y<1 < ,Y < !) (a) Find P (X 1 2 (b) Find P(X < 2Y)
Answer #1

a)P(X<1/2 ,Y<1/4)=\int_{0}^{1/2}\int_{0}^{1/4} (x+y) dy dx =\int_{0}^{1/2}(xy+y2/2) |1/40 dx =\int_{0}^{1/2} (x/4+1/32) dx =(x2/8+x/32) |1/20

=(1/32+1/64) =3/64

b)P(X<2Y) =1-P(Y<X/2) =1 -Jo\int_{0}^{x/2} (x+y) dy dx =1-Jo (xy+y2/2)|x/20 dx =1-Jo (x2/2+x2/8) dx

=1- Jo 5x2/8 dx =1-(5x3/24)|10 =1-5/24 =19/24